OpenStudy (habbababba1):

will fan and medal!!!!.....If a line crosses the y-axis at (0, 1) and has a slope of 4/5, what is the equation of the line?

OpenStudy (igreen):

Which form are you looking for? Slope-intercept?

OpenStudy (habbababba1):

im not sure i just know im looking for the equation that it would make

OpenStudy (habbababba1):

i have answer choices

OpenStudy (habbababba1):

4y − 5x = 5 y − 4x = 5 5y + 4x = 5 5y − 4x = 5

OpenStudy (igreen):

That's standard form. First put it in slope-intercept form. Do you know how? y = mx + b Where 'm' is the slope and 'b' is the y-intercept.

OpenStudy (igreen):

The question already gave you the slope and the y-intercept, all you have to do is plug them in and convert to standard form.